# Deep-ELA: Deep Exploratory Landscape Analysis with Self-Supervised Pretrained Transformers for Single- and Multi-Objective Continuous Optimization Problems

### v1

@@ 5.2 High-Level Property Prediction
Table 3: F1-Score with macro aggregation (F1-Scores a computed individually per class and then mean-aggregated over all classes) of predicting High-Level Properties of all 24 BBOB functions and instances $126$ to $150$. The results marked with a $*{}^{*}$ are directly taken from Seiler et al., (2022) while the ones marked with ${}^{**}$ are taken from van Stein et al., (2023). The last eight columns list the performances obtained by our proposed Deep-ELA models. van Stein et al., (2023) did not provide results for $d=3$ and aggregated over all dimensions (all). Our medium models cannot handle data with more than six total dimensions. Therefore, there are no results for $d=10$ for the medium models.
@@ 5.3 Single-Objective Automated Algorithm Selection
Table 4: Comparison of the relative ERT (relERT) values for the algorithm selection study on instances $1$ to $5$ of the 24 single-objective BBOB functions. The first five columns (marked with $*{}^{*}$) are directly taken from Prager et al., (2022), while the last eight show the performance of our proposed Deep-ELA approach. Our medium models cannot handle data with more than six dimensions. Therefore, there are no results for $d=10$ for the medium models.

### v2

@@ 5.2 High-Level Property Prediction
Table 3: F1-Score with macro aggregation (F1-Scores a computed individually per class and then mean-aggregated over all classes) of predicting High-Level Properties of all 24 BBOB functions and instances $126$ to $150$. The results marked with a are directly taken from Seiler et al., (2022) while the ones marked with ∗∗ are taken from van Stein et al., (2023). The last eight columns list the performances obtained by our proposed Deep-ELA models. van Stein et al., (2023) did not provide results for $d=3$ and aggregated over all dimensions (all). Our medium models cannot handle data with more than six total dimensions. Therefore, there are no results for $d=10$ for the medium models.
@@ 5.3 Single-Objective Automated Algorithm Selection
Table 4: Comparison of the relative ERT (relERT) values for the algorithm selection study on instances $1$ to $5$ of the 24 single-objective BBOB functions. The first five columns (marked with ) are directly taken from Prager et al., (2022), while the last eight show the performance of our proposed Deep-ELA approach. Our medium models cannot handle data with more than six dimensions. Therefore, there are no results for $d=10$ for the medium models.